# A Spectral Localization Method for Time-Fractional Integro-Differential Equations with Nonsmooth Data

### v1

@@ Abstract
Keywords Time-fractional integro-differential equation  $\cdot$ Nonsmooth/vanishing data  $\cdot$ Contour integral method  $\cdot$ Localization scheme  $\cdot$ Error estimates
@@ 1 Introduction
$\begin{cases}{}_{0}^{C}D_{t}^{\beta}u+Au+\int_{0}^{t}\kappa_{\alpha}(t-s)Au(s)\,ds=f,&(x,t)\in\Omega\times(0,T],\\ u=0,&(x,t)\in\partial\Omega\times(0,T],\\ u(0)=u_{0}(x),&x\in\Omega,\end{cases}$ (1.1)
where $^{C}_{0}D_{t}^{\beta},\beta\in(0,1),$ denotes the Caputo fractional derivative, given as
$^{C}_{0}D_{t}^{\beta}u(x,t):=\frac{1}{\Gamma(1-\beta)}\int_{0}^{t}\frac{\partial u(x,\tau)}{\partial\tau}(t-\tau)^{-\beta}\,d\tau,$ (1.2)
$A:=-\sum_{i=1}^{d}\frac{\partial^{2}}{\partial x_{i}^{2}}$ represents the negative Laplace operator, $\kappa_{\alpha}(t),\alpha\in(0,1),$ is a memory kernel function defined by
@@
with Gamma function $\Gamma(s)=\int_{0}^{\infty}t^{s-1}e^{-t}\,dt,$ the function $f=f(x,t)$ is the source term, $\Omega$ is a bounded convex polygonal domain in $R^{n}(n\geq 1)$ with boundary $\partial\Omega$, $T$ is a fixed positive number, and $u_{0}(x)$ is the initial data.
This type of equation, with the nonlocal operator $^{C}_{0}D_{t}^{\beta}$, depicts the anomalous motion characteristics of particles in sub-diffusion process (where the mean squared displacement follows a power law $\langle x^{2}(t)\rangle\propto t^{\beta}$ with $\beta\in(0,1)$), and can better describe the physical phenomena of anomalous diffusion, including the transport of proteins within the cytoskeleton and the diffusion of pollutants in porous media [6]. Additionally, the memory kernel function $\kappa_{\alpha}(t)$ demonstrates a power-law decay characteristics, effectively models the long-time scale dynamics of processes such as the relaxation response in electrochemical systems [29]. Moreover, from a microscopic perspective, $u(x,t)$ describes the probability density of particles at time $t$ and position $x$, while at the macroscopic scale, it corresponds to the material concentration, and $f(x,t)$ represents the external force that is artificially applied or given. Such integro-differential equation can better describe physical processes that exhibit anomalous diffusion characteristics and long-range temporal dependence, such as telomere motion in cells [26], diffusion in porous catalysts [3], and transport in nuclear reactor materials [39]. Owing to its certain application potential, problem (1.1) and its nonlinear variants have attracted considerable research attention, with wide-ranging studies spanning both theoretical analysis andwide-range numerical computation.
For theoretical analysis, the authors in [11] study the existence and regularity of the solutions to problem (1.1) in Sobolev spaces with Riemann–Liouville fractional derivatives. Regularity estimates for both homogeneous and nonhomogeneous versions of equation (1.1) are established in [22] under a range of regularity assumptions on the initial data and source term, with particular emphasis on nonsmooth initial data; corresponding stability estimates are also provided. The authors of [9] investigate the semilinear variant of (1.1) in one dimension, establishing the existence and regularity of solutions within fractional Hölder spaces. These results are later extended to the multidimensional case in [10], which also provides a further analysis of the well-posedness of the solution. Their contributions plays a key role in advancing the theoretical understanding of such equations, leading to a relatively well-developed foundation for the analysis.
@@
where $\gamma\geq 0$ and $D_{t}^{\beta}$ is the Riemann-Liouville or Caputo fractional derivative, the authors in [25] apply a compact difference scheme to solve (1.3) with a particular weakly singular kernel for $\alpha$= 1/2, demonstrating that the method achieves an accuracy of $O(\tau+h^{4})$. Based on this, a compact finite difference scheme is proposed in [19] that enhances the temporal convergence order to $O(\tau^{2}+h^{4})$ by using of the weighted and shifted Grünwald formula and compact difference operators. Moreover, the authors in [40] develop an Alternating Direction Implicit (ADI) difference scheme for solving (1.3), where the Caputo fractional derivative is discretized using the L1 formula, while the integral term is approximated via the convolution quadrature. The proposed scheme is shown to achieve a convergence rate of $O(\tau^{\min\{2-\beta,1+\alpha\}}+h^{2})$. Then a compact L1-ADI scheme is developed in [35], which improves the temporal convergence by employing the trapezoidal Proportional-Integral (PI) rule. In addition, the authors of [28] and [24] investigate the special case of (1.3) with $\beta=1$ and $\gamma=0$. Under suitable assumptions on the data and for convex or sufficiently smooth spatial domains, they demonstrate that the error converges at the rate of $O(\tau^{2}+h^{2})$.
@@
In the aforementioned studies, the proposed algorithms are inherently nonlocal, leading to a substantial increase in both computational complexity and storage demands as the problem scale expands. Moreover, most of these methods achieve at most second-order convergence in time. Another limitation is their reliance on high solution regularity—for instance, $u(t)\in C^{2}(\Omega)$ is required in [25]—an assumption often inconsistent with real-world applications. In practice, nonsmooth initial data are frequently encountered, and the actual regularity of solutions tends to be limited. To overcome these challenges, the authors in [22] present an analysis of the L1 scheme applied to problem (1.1), deriving optimal error estimates that remain valid even for nonsmooth initial data such as $u_{0}\in L^{2}(\Omega)$.
@@
Over the past two decades, a technique known as the Contour Integral Method (CIM), which is based on Laplace transforms and contour integration, has gained prominence [18, 7, 13]. The fundamental idea behind this approach can be traced back to 1979, when Talbot provided a systematic account of its origins and development [34]. This method expresses the solution of the differential equation via an inverse Laplace transform, which is then approximated by numerical integration along a suitably chosen contour. The idea was first applied to parabolic equations by [31], and later extended to fractional-order evolution equations in [23]. Subsequently, the CIM has been extended to fractional differential equations. For instance, [12] introduces a parallel Laplace transform-based approach for backward parabolic equations; [7] developes an improved CIM for solving the Black-Scholes and Heston equations; a parabolic contour CIM is applied in [16] to handle nonlinear subdiffusion equations with nonsmooth initial data. Collectively, these studies demonstrate that, compared with traditional numerical methods, CIM offers several key advantages: (1) it accommodates solutions with low regularity; (2) the method is localized, meaning the solution at each time level can be computed independently without historical data, which facilitates efficient parallelization; (3) it achieves spectral accuracy in time.
@@
The aim of this paper is to develop a high-performance numerical algorithm—a localizition method—for the nonlocal problem (1.1). This method is designed to handle solutions with low regularity while achieving high accuracy, thereby addressing challenges related to nonlocality and nonsmoothness. Specifically, for temporal discretization, we employ the contour integral method (CIM), which represents the solution via an inverse Laplace transform expressed as a contour integral along the left branch of a parameterized hyperbolic contour. This integral is then approximated using the midpoint rule, leading to a semi-discrete CIM scheme in time. For spatial discretization, the standard Galerkin finite element method (FEM) is applied. Additionally, we conduct a rigorous error analysis of the fully discrete scheme; the main theoretical results are summarized in Theorem 3 and Theorem 4.
@@
  • A novel numerical scheme is developed, which uses a parameterized hyperbolic contour CIM for temporal discretization to localize the originally nonlocal problem.
  • The proposed algorithm imposes very low regularity requirements on the solution, enabling robust performance even in the presence of nonsmooth initial data.
  • Rigorous error analysis demonstrates that the method achieves second-order convergence in space and spectral accuracy in time.
@@
In this section, we introduce the fundamental concepts of the CIM and employ it for semi-discretization in time. Spatial semi-discretization is then carried out using the standard Galerkin FEM. Building on these semi-discrete formulations, we present a fully discrete scheme for problem (1.1) and conclude the section with an accelerated algorithm for its efficient implementation.
@@ 2.1 Contour Integral Method (CIM)
Let $u(t)$ be a real-valued function. Under certain conditions-specifically, if there exist constants $\sigma_{0}>0$ and $M>0$ such that $|u(t)|\leq Me^{\sigma_{0}t}$-then its Laplace transform $\widehat{u}(z):=\mathcal{L}\{u(t)\}(z)$ exists for $\operatorname{Re}(z)>\sigma_{0}$. In that case, $u(t)$ can be expressed via the inverse Laplace transform as
$u(t)=\frac{1}{2\pi i}\int_{\sigma_{0}-i\infty}^{\sigma_{0}+i\infty}e^{zt}\widehat{u}(z)\,dz,\quad Re(z)>\sigma_{0}.$ (2.1)
To compute $u(t)$, it then suffices to choose a suitable numerical method for approximating this contour integral. In order to avoid the high-frequency oscillations of the exponential factor along the vertical line $z=\sigma_{0}+iy$, where $e^{zt}=e^{\sigma_{0}t}\cdot\left(\cos(yt)+i\sin(yt)\right)$ for $\infty<y<+\infty$, we employ the CIM, which offers both efficiency and simplicity.
@@
Figure 1: Illustration of the integration path transformation
@@
The integral (2.3) can then be efficiently approximated using either the trapezoidal rule or the midpoint rule.
@@
$u(t)$ $\approx I_{\tau}=\frac{\tau}{2\pi i}\sum_{k=-\infty}^{+\infty}e^{z(\phi_{k})t}\widehat{u}(z(\phi_{k}))z^{\prime}(\phi_{k})$ (2.5)
$\approx I_{\tau,N}=\frac{\tau}{2\pi i}\sum_{k=1-N}^{N-1}e^{z(\phi_{k})t}\widehat{u}(z(\phi_{k}))z^{\prime}(\phi_{k}),$
where $\tau$ is a uniform time step size along the $\phi$-axis, NN denotes the number of quadrature nodes, and $\phi_{k}:=\left(k+\frac{1}{2}\right)\tau,k=0,\pm 1,\ldots,\pm(N-1)$.
The scheme (2.5) is referred to as the temporal CIM scheme for approximating $u(t)$, which attains spectral accuracy under appropriate conditions. As shown in [37], the CIM imposes low regularity requirements on $u(t)$-it only necessitates the existence of the Laplace transform. In fact, the convergence behavior of (2.5) is closely linked to the analyticity of $\widehat{u}(z)$: the width of the region of analytic continuation of $\widehat{u}(z)$ around the contour $\Gamma$ is positively correlated with the exponential convergence rate of the CIM error.
@@ 2.2 Temporal Semi-discrete CIM Scheme
According to the key idea of the CIM, we begin by deriving the Laplace transform of $u(x,t)$, hereafter abbreviated as $u(t)$, from problem (1.1). Since the Laplace transform of the Caputo fractional derivative (1.2) with $\beta\in(0,1)$ is given by
$\mathcal{L}\left\{{}_{0}^{C}{D}_{t}^{\beta}u(t)\right\}=z^{\beta}\widehat{u}(z)-z^{\beta-1}u(0),$
@@
$\mathcal{L}\left\{\int_{0}^{t}\kappa_{\alpha}(t-s)Au(s)\,ds\right\}=\frac{1}{z^{\alpha}}A\widehat{u}(z),$
@@
$\widehat{u}(z)=\left(m(z)I+A\right)^{-1}\left(\frac{m(z)}{z}u_{0}+\frac{m(z)}{z^{\beta}}\widehat{f}(z)\right),$ (2.6)
@@
By appropriately choosing the range of values for $z$, it is possible to ensure that $\left(m(z)I+A\right)^{-1}$ remains free of singularities. A detailed discussion of this will be provided in Section 3. So, now one can observe that the singularity of $\widehat{u}(z)$ only occurs at $z=0$, since $z^{\alpha}=-1$ does not hold when $z$ is within a single-valued branch.
@@
$\Gamma:z(\phi)=\mu\left(1+\sin(i\phi-\theta)\right),\quad\phi=\nu+ic\in K,\quad-\infty<\nu<\infty,$ (2.8)
@@
$K=\left\{\phi=\nu+ic\in\mathbb{C}\,\middle|\,\nu\in\mathbb{R},\ |c|<\tilde{c}\right\}$,
with $\tilde{c}$ to be determined. One reason for choosing a hyperbolic contour is that, as observed in [37], such contours generally require fewer discretization points $N$ than other previously mentioned types to achieve the same level of accuracy. Another reason is that hyperbolas possess asymptotes, which are necessary for determining the optimal value of $\tilde{c}$. Further details on this matter will be provided in Section 3.
To facilitate analysis and discussion, we express Eq. (2.8) in hyperbolic form:
$\left(\frac{a-\mu}{\sin(\theta+c)}\right)^{2}-\left(\frac{b}{\cos(\theta+c)}\right)^{2}=\mu^{2},\quad z=a+ib.$ (2.9)
Figure 2 provides a visual representation of the mapping between the horizontal region $K$ in the $\phi$-plane and its corresponding region $S$ in the $z$-plane. Actually, from Eq. (2.9), we observe that when $c=0$, the contour $\Gamma$ corresponds to the left branch of a hyperbola, shown in white on the right side of Figure 2. As $c$ decreases from 0 (i.e., $c<0$), indicating a downward shift of the horizontal line in region $K$, the corresponding hyperbola in region $S$ expands and shifts rightward. At the minimum value $c=-\theta$, its image in the $z$-plane becomes a vertical line. Conversely, when $c>0$ and increases from 0, corresponding to an upward movement of the line in the $\phi$-plane, the hyperbola gradually closes and shifts leftward, eventually coinciding with the negative real axis when $c$ reaches $\pi/2-\theta$.
Figure 2: Mapping relationship between the horizontal region $K$ in the $\phi$-plane and the region $S$ in the $z$-plane
@@
$u^{N}(t):=I_{\tau;N}=\frac{\tau}{\pi}\operatorname{Im}\left\{\sum_{k=0}^{N-1}e^{z(\phi_{k})t}\widehat{u}(z(\phi_{k}))z^{\prime}(\phi_{k})\right\},$ (2.10)
where $\phi_{k}=\left(k+\frac{1}{2}\right)\tau$ for $k=0,1,\ldots,N-1$. Here, $\hat{u}(z(\phi_{k}))$ and $z(\phi_{k})$ denote the values of (2.6) and (2.8) at $\phi=\phi_{k}$, respectively. One can observe that compared to (2.5), the scheme (2.10) reduces the computational cost by half. For notational convenience, we denote $z(\phi_{k})$ as $z_{k}$ in the subsequent analysis.

### v2

@@ Abstract
Keywords: Time-fractional integro-differential equation; Nonsmooth/vanishing data; Contour integral method; Localization scheme; Error estimates
AMS subject classifications: 65R10, 65M15, 65M60, 35R09, 35R05
@@ 1 Introduction
$\left\{\begin{aligned} &{{}_{0}^{C}}\!D_{t}^{\beta}u+Au+\int_{0}^{t}\kappa_{\alpha}(t-s)Au(s)\,ds=f,&&(x,t)\in\Omega\times(0,T],\\ &u(x,t)=0,&&(x,t)\in\partial\Omega\times(0,T],\\ &u(x,0)=u_{0}(x),&&x\in\Omega,\end{aligned}\right$. (1.1)
where ${{}_{0}^{C}}\!D_{t}^{\beta}$ with $\beta\in(0,1)$ denotes the Caputo fractional derivative given by
${{}_{0}^{C}}\!D_{t}^{\beta}u(x,t):=\frac{1}{\Gamma(1-\beta)}\int_{0}^{t}\frac{\partial u(x,\tau)}{\partial\tau}(t-\tau)^{-\beta}\,d\tau,$ (1.2)
$\kappa_{\alpha}(t)$ with $\alpha\in(0,1)$ is a memory kernel function defined by
@@
and $\Gamma(s)=\int_{0}^{\infty}t^{s-1}e^{-t}\,dt$ stands for the Euler Gamma function. In (1.1), $\Omega$ is a bounded convex polygonal domain in $\mathbb{R}^{n}(n\geq 1)$ with boundary $\partial\Omega$, $T>0$ is a fixed final time, $A:=-\sum_{i=1}^{d}\frac{\partial^{2}}{\partial x_{i}^{2}}$ represents the negative Laplace operator, the function $f=f(x,t)$ is the source term, and $u_{0}(x)$ is the initial data.
This type of equation, with the nonlocal operator ${{}_{0}^{C}}\!D_{t}^{\beta}$, depicts the anomalous motion characteristics of particles in sub-diffusion process (where the mean squared displacement follows a power law $\langle x^{2}(t)\rangle\propto t^{\beta}$ with $\beta\in(0,1)$), and can better describe the physical phenomena of anomalous diffusion, including the transport of proteins within the cytoskeleton and the diffusion of pollutants in porous media [7]. Additionally, the memory kernel function $\kappa_{\alpha}(t)$ demonstrates a power-law decay characteristics, effectively models the long-time scale dynamics of processes such as the relaxation response in electrochemical systems [32]. Moreover, from a microscopic perspective, $u(x,t)$ describes the probability density of particles at time $t$ and position $x$, while at the macroscopic scale, it corresponds to the material concentration, and $f(x,t)$ represents the external force that is artificially applied or given. Such integro-differential equation can better describe physical processes that exhibit anomalous diffusion characteristics and long-range temporal dependence, such as telomere motion in cells [28], diffusion in porous catalysts [4], and transport in nuclear reactor materials [41]. Owing to its certain application potential, problem (1.1) and its nonlinear variants have attracted considerable research attention, with wide-ranging studies spanning both theoretical analysis and numerical computation.
For theoretical analysis, the authors in [1] studied the existence and regularity of the solutions to problem (1.1) in Sobolev spaces with Riemann-Liouville fractional derivatives. Regularity estimates for both homogeneous and nonhomogeneous versions of equation (1.1) were established in [22] under a range of regularity assumptions on the initial data and source term, with particular emphasis on nonsmooth initial data; corresponding stability estimates were also provided. The authors of [10] investigated the semilinear variant of (1.1) in one dimension, establishing the existence and regularity of solutions within fractional Hölder spaces. These results were later extended to the multidimensional case in [11], which also provided a further analysis of the well-posedness of the solution. Their contributions play a key role in advancing the theoretical understanding of such equations, leading to a relatively well-developed foundation for the analysis.
@@
where $\gamma\geq 0$ and $D_{t}^{\beta}$ is the Riemann-Liouville or Caputo fractional derivative, the authors in [27] applied a compact finite difference scheme to solve (1.3) with a particular weakly singular kernel for $\alpha=1/2$, demonstrating that the method achieves an accuracy of $O(\tau+h^{4})$. Based on this, a compact finite difference scheme was proposed in [19] that enhances the temporal convergence order to $O(\tau^{2}+h^{4})$ by using the weighted and shifted Grünwald formula and compact difference operators. Moreover, the authors in [42] developed an Alternating Direction Implicit (ADI) difference scheme for solving (1.3), where the Caputo fractional derivative is discretized by the L1 formula, while the integral term is approximated via the convolution quadrature. The proposed scheme is shown to achieve a convergence rate of $O(\tau^{\min\{2-\beta,1+\alpha\}}+h^{2})$. Then a compact L1-ADI scheme was developed in [37], which improves the temporal convergence by employing the trapezoidal Proportional-Integral (PI) rule. In addition, the authors of [30] and [26] investigated the special case of (1.3) with $\beta=1$ and $\gamma=0$. Under suitable assumptions on the data and for convex or sufficiently smooth spatial domains, they demonstrated that the error converges at the rate of $O(\tau^{2}+h^{2})$.
@@
In the aforementioned studies, the proposed algorithms are inherently nonlocal, leading to a substantial increase in both computational complexity and storage demands as the problem scale expands. Moreover, most of these methods achieve at most second-order convergence in time. Another limitation is their reliance on high solution regularity – for instance, $u(t)\in C^{2}[0,T]$ is required in [42] – an assumption often inconsistent with real-world applications. In practice, nonsmooth initial data are frequently encountered, and the actual regularity of solutions tends to be limited. To overcome these challenges, the authors in [22] presented an analysis of the L1 scheme applied to problem (1.1), deriving optimal error estimates that remain valid even for nonsmooth initial data such as $u_{0}\in L^{2}(\Omega)$.
@@
Over the past two decades, a technique known as the Contour Integral Method (CIM), which is based on Laplace transforms and contour integration, has gained prominence [18, 8, 13]. The fundamental idea behind this approach can be traced back to 1979, when Talbot provided a systematic account of its origins and development [36]. This method expresses the solution of the differential equation via an inverse Laplace transform, which is then approximated by numerical integration along a suitably chosen contour. The idea was first applied to parabolic equations by [33], and later extended to fractional-order evolution equations in [25]. Subsequently, the CIM has been extended to fractional differential equations. For instance, [12] introduced a parallel Laplace transform-based approach for backward parabolic equations; [8] developd an improved CIM for solving the Black-Scholes and Heston equations; a parabolic contour CIM was applied in [16] to handle nonlinear subdiffusion equations with nonsmooth initial data. Collectively, these studies demonstrate that, compared with traditional numerical methods, CIM offers several key advantages: (1) it accommodates solutions with low regularity; (2) the method is localized, meaning the solution at each time level can be computed independently without historical data, which facilitates efficient parallelization; (3) it achieves spectral accuracy in time.
@@
The aim of this paper is to develop a high-performance numerical algorithm – a localization method – for the nonlocal problem (1.1). This method is designed to handle solutions with low regularity while achieving high accuracy, thereby addressing challenges related to nonlocality and nonsmoothness. Specifically, for temporal discretization, we employ the contour integral method (CIM), which represents the solution via an inverse Laplace transform expressed as a contour integral along the left branch of a parameterized hyperbolic contour. This integral is then approximated using the midpoint rule, leading to a semi-discrete CIM scheme in time. For spatial discretization, the piecewise linear finite element method (FEM) is applied. Additionally, we conduct a rigorous error analysis of the fully discrete scheme; the main theoretical results are summarized in Theorems 3 and 4.
@@
  • An efficient numerical scheme is developed, which uses a parameterized hyperbolic contour CIM for temporal discretization to localize the originally nonlocal problem.
  • The proposed algorithm imposes very low regularity requirements on the solution, enabling robust performance even in the presence of nonsmooth initial data and low regularity solution.
  • Rigorous error analysis demonstrates that the method achieves spectral accuracy in time and second-order convergence in space.
@@
In this section, we introduce the fundamental concepts of the CIM and employ it for semi-discretization in time. Spatial semi-discretization is then carried out using the piecewise linear FEM. Building on these semi-discrete formulations, we present a fully discrete scheme for problem (1.1) and conclude the section with an accelerated algorithm for its efficient implementation.
@@ 2.1 Contour Integral Method (CIM)
Let $u(t)$ be a real-valued function. Under certain conditions-specifically, if there exist constants $\sigma_{0}>0$ and $M>0$ such that $|u(t)|\leq Me^{\sigma_{0}t}$, then its Laplace transform $\widehat{u}(z):=\mathcal{L}\{u(t)\}(z)$ exists for $\operatorname{Re}(z)>\sigma_{0}$. In that case, $u(t)$ can be expressed via the inverse Laplace transform as
$u(t)=\frac{1}{2\pi i}\int_{\sigma_{0}-i\infty}^{\sigma_{0}+i\infty}e^{zt}\widehat{u}(z)\,dz,\quad\operatorname{Re}(z)>\sigma_{0}.$ (2.1)
To compute $u(t)$, it then suffices to choose a suitable numerical method for approximating this contour integral. In order to avoid the high-frequency oscillations of the exponential factor along the vertical line $z=\sigma_{0}+iy$, where $e^{zt}=e^{\sigma_{0}t}\cdot(\cos(yt)+i\sin(yt))$ for $\infty<y<+\infty$, we employ the CIM, which offers both efficiency and simplicity.
@@
Figure 1: Illustration of the integration path transformation.
@@
The integral (2.3) can then be efficiently approximated using either the trapezoidal rule or the midpoint rule [23, 24].
@@
$u(t)$ $\approx I_{\tau}=\tau\sum_{k=-\infty}^{+\infty}g(\phi_{k},t)\approx I_{\tau,N}=\tau\sum_{k=1-N}^{N-1}g(\phi_{k},t),$ (2.5)
where $\tau$ is a uniform time step size along the $\phi$-axis, $N$ denotes the number of quadrature nodes, and $\phi_{k}:=(k+\frac{1}{2})\tau,k=0,\pm 1,\ldots,\pm(N-1)$.
The scheme (2.5) is referred to the temporal CIM scheme for approximating $u(t)$, which attains spectral accuracy under appropriate conditions. As shown in [39], the CIM imposes low regularity requirements on $u(t)$, and it only necessitates the existence of the Laplace transform. In fact, the convergence behavior of (2.5) is closely linked to the analyticity of $\widehat{u}(z)$: the width of the region of analytic continuation of $\widehat{u}(z)$ around the contour $\Gamma$ is positively correlated with the exponential convergence rate of the CIM error.
The error of the CIM scheme (2.5) can be decomposed into the discretization error $DE:=\|I-I_{\tau}\|_{L^{2}(\Omega)}$ and truncation error $TE:=\|I_{\tau}-I_{\tau;N}\|_{L^{2}(\Omega)}$,
$\|I-I_{\tau,N}\|_{L^{2}(\Omega)}\leq\|I-I_{\tau}\|_{L^{2}(\Omega)}+\|I_{\tau}-I_{\tau;N}\|_{L^{2}(\Omega)}=:DE+TE,$
where the definitions of $I$, $I_{\tau}$ and $I_{\tau;N}$ are given by (2.3) and (2.5), respectively. In [39], the authors analyzed these two types of errors by considering the function $g(\phi,t)$ in (2.3) as complex-valued, i.e., $g(\nu+ic,t)$, where $g(\phi,t)$ admits an analytic continuation w.r.t. $\phi=\nu+ic$, and the time $t$ can be regarded as a given parameter. They assume that
[…]
which leads to the following estimate for the discretization error:
$DE\leq\frac{M(\tilde{c},t)}{e^{2\pi\tilde{c}/\tau}-1},$ (2.6)
implying that $DE=O(e^{-2\pi\tilde{c}/\tau})$. Estimate (2.6) indicates that larger values of $\tilde{c}$, which represents the width of the region of analytic continuation of $\widehat{u}(z(\phi))$, result in smaller discretization errors and faster convergence. As for the truncation error $TE$, assuming that $g(\phi,t)$ decays rapidly as $\phi=\nu+ic$, with $\nu\to\pm\infty$, it can be approximated by the magnitude of the last retained term. That is, for fixed $\tau$ and as $N\to\infty$,
$TE=O(|g(\tau N,t)|).$ (2.7)
@@ 2.2 Temporal Semi-discrete CIM Scheme
According to the key idea of the CIM, we begin by deriving the Laplace transform of $u(x,t)$, hereafter abbreviated as $u(t)$, from problem (1.1). Since the Laplace transform of the Caputo fractional derivative (1.2) with $\beta\in(0,1)$ is given by [31]
$\mathcal{L}\big\{{{}_{0}^{C}}\!D_{t}^{\beta}u(t)\big\}=z^{\beta}\widehat{u}(z)-z^{\beta-1}u(0),$
@@
$\mathcal{L}\Big\{\int_{0}^{t}\kappa_{\alpha}(t-s)Au(s)\,ds\Big\}=\frac{1}{z^{\alpha}}A\widehat{u}(z),$
@@
$\widehat{u}(z)=\big(m(z)I+A\big)^{-1}\Big(\frac{m(z)}{z}u_{0}+\frac{m(z)}{z^{\beta}}\widehat{f}(z)\Big),$ (2.8)
@@
By appropriately choosing the range of values for $z$, it is possible to ensure that $(m(z)I+A)^{-1}$ remains free of singularities. A detailed discussion will be provided in Section 3. We can then observe that the singularity of $\widehat{u}(z)$ only occurs at $z=0$, since $z^{\alpha}=-1$ does not hold when $z$ is within a single-valued branch.
@@
$\Gamma:z(\phi)=\mu\big(1+\sin(i\phi-\theta)\big),\quad\phi=\nu+ic\in K,\quad-\infty<\nu<\infty,$ (2.10)
@@
$K=\big\{\phi=\nu+ic\in\mathbb{C}~|~\nu\in\mathbb{R},~|c|<\tilde{c}\big\}$,
with $\tilde{c}$ to be determined. One reason for choosing a hyperbolic contour is that, as observed in [39], such contours generally require fewer discretization points $N$ than other previously mentioned types to achieve the same level of accuracy. Another reason is that hyperbolas possess asymptotes, which are necessary for determining the optimal value of $\tilde{c}$. Further details on this matter will also be provided in Section 3.
To facilitate analysis and discussion, we express (2.10) in hyperbolic form:
$\Big(\frac{a-\mu}{\sin(\theta+c)}\Big)^{2}-\Big(\frac{b}{\cos(\theta+c)}\Big)^{2}=\mu^{2},\quad z=a+ib.$ (2.11)
Figure 2 provides a visual representation of the mapping between the horizontal region $K$ in the $\phi$-plane and its corresponding region $S$ in the $z$-plane. Actually, from (2.11), we observe that when $c=0$, the contour $\Gamma$ corresponds to the left branch of a hyperbola, shown in white on the right side of Figure 2. As $c$ decreases from 0 (i.e., $c<0$), indicating a downward shift of the horizontal line in region $K$, the corresponding hyperbola in region $S$ expands and shifts rightward. At the minimum value $c=-\theta$, its image in the $z$-plane becomes a vertical line. Conversely, when $c>0$ and increases from 0, corresponding to an upward movement of the line in the $\phi$-plane, the hyperbola in the $z$-plane gradually closes and shifts leftward, eventually coinciding with the negative real axis when $c$ reaches $\pi/2-\theta$.
Figure 2: A schematic mapping relationship between the horizontal region $K$ in the $\phi$-plane and the region $S$ in the $z$-plane.
@@
$u^{N}(t):=I_{\tau;N}=\frac{\tau}{\pi}\operatorname{Im}\Big\{\sum_{k=0}^{N-1}e^{z(\phi_{k})t}\widehat{u}(z(\phi_{k}))z^{\prime}(\phi_{k})\Big\},$ (2.12)
where $\phi_{k}:=(k+\frac{1}{2})\tau$ for $k=0,1,\ldots,N-1$. Here, $\widehat{u}(z(\phi_{k}))$ and $z(\phi_{k})$ denote the values of (2.8) and (2.10) at $\phi=\phi_{k}$, respectively. One can observe that compared to (2.5), the scheme (2.12) reduces the computational cost by half. For notational convenience, we denote $z(\phi_{k})$ as $z_{k}$ in the subsequent analysis.